# SUBA-Jet: a new Model for Jets in Heavy Ion Collisions

### v1

@@ III.2 Medium modifications
$\hat{q}_{\rm GA}(T,p)=\hat{q}_{\rm GA}(T,p=20\text{ GeV})\times\hat{q}_{\rm cof}(p)\,\quad\text{where}\quad\hat{q}_{\rm cof}(p)=\frac{1.69+1.25\,p}{4.07+p+0.85\ln\left(p+1\right)}\,,$ (16)
satisfying $\hat{q}_{\rm cof}(20\text{ GeV})=1$ (dimensionless). In this definition, $p$ is the dimensionless numerical value of the momentum expressed in GeV. For gluons, a color factor of $C_{A}/C_{F}$ is applied.
@@
The consequences of the in-medium induced virtuality increase is shown in Figs. 2 and 3 for initial parton energies of $10$ and $50$ GeV. We see an increase of the virtuality, which starts after the partons have travelled a mean free path and which may reach $50\%$ for the sum of the squared virtualities shown on Fig. 3. A higher virtuality leads to more splittings in which the virtuality is reduced, as shown in Fig. 4. In Fig. 5 the fraction of the initial tagged partons, which have reached the minimum virtuality of $2Q_{0}$, is shown for different initial energies as a function of time. The consequences of the presence of a medium (eq. (14) ) are also shown. The medium-induced virtuality increase has the consequence that the initial quark reaches the minimum virtuality earlier. For larger jet energies, the effect of the medium modifications plays a bigger role because the virtuality increase is larger and therefore the loss of virtuality proceeds faster. For an initial energy $E_{\text{ini}}=10$ GeV in about $80\%$ of the jets the initial quark has reached after around $1$ fm$/c$ the minimal virtuality and has entered the low virtuality regime. For $E_{\text{ini}}=50$ GeV this happens after about $2$ fm$/c$.
@@ IV.1 Thermal masses
$m_{q}^{\text{therm}}(T)=0.087\text{ GeV}+0.7\times T\,.$ (21)
In the code we use these linear fit functions.
@@ IV.2 Coupling constant
During the evolution of the QGP, the elastic collisions will be probed especially at small momentum transfer. For this reason, we neglect the dependence of $\alpha_{s}$ on $Q^{2}$ and consider only an effective temperature dependence of $\alpha_{s}$ in eqs. (25) to (28). Our method relies on the models developed in ref. Gossiaux:2008jv that have led to a good agreement with experimental data in the HF sector. This is in particular true for the running model (E). We have evaluated the $\hat{q}(T,p)$ for model (E) — adapted to the light-flavor case — and then fitted $\alpha_{s}(T)$ of model (C) to reproduce at best these $\hat{q}$ values. This method leads to
@@ IV.3 Elastic collisions
Thanks to the IR regulator, the total elastic collision rate does not diverge, however its second moment - the transport coefficient $\hat{q}=\int q_{T}^{2}\frac{\mathrm{d}\Gamma_{\rm el}}{\mathrm{d}^{2}q_{T}}/\int\frac{\mathrm{d}\Gamma_{\rm el}}{\mathrm{d}^{2}q_{T}}$ does if $q_{T}$ is not limited from above. Therefore, to match the effective $\hat{q}$ of a more exact approach, model C of ref. Gossiaux:2008jv , which uses HTL propagators, we limit the transverse momentum exchange to $q_{T}^{2}<q_{T,{\rm max}}^{2}=2ET$, where $E$ is the energy of the incoming parton [^9] Not to be confused with the initial jet parton. With this upper boundary one obtains a finite transport coefficient for jet quarks:
@@ IV.4 Inelastic collisions
The elementary inelastic collision, which we employ in our approach, is the radiation of a virtual gluon from a jet parton. The formalism, which we apply, is the QCD extension of the QED Bremsstrahlung and has been studied by Gunion and Bertsch (GB) Gunion:1981qs . As discussed, the emitted gluon with a four-vector $(\omega,k_{{\perp}},k_{z})$ needs a time $t_{f}=\omega/k_{{\perp}}^{2}$, [^10] In the numerical code, gluons are also radiated from secondary partons; in this case, the notation $\perp$” is slightly abusive, although still pretty clear. called formation time, until it becomes real. If the formation time is shorter than the mean free path $\lambda$ of the incoming parton in the medium the formation time is shorter than the mean free path between two collisions, gluon emission from the different vertices is independent and we are in the Bethe Heitler (BH) regime. If, however, the formation time is larger than the mean free path, which depends on the temperature via the density ($\tau_{f}\propto T$), the gluon emission amplitudes interfere destructively, a process which has been studied in the BDMPS-Z approach Baier:2001yt ; Baier:1998yf ; Baier:1996sk . In this regime the gluon emission $\omega\mathrm{d}N/\mathrm{d}\omega$ is suppressed with respect to the BH regime. If the formation length exceeds the length $L$ of the QGP (assumed to be represented by a box of length L), which is the case if $\omega$ exceeds the critical value $\omega_{c}\sim\hat{q}L^{2}$, gluons will, on the average, not be formed anymore and will only be emitted by fluctuations. The three $\omega$ regions are schematically shown in Fig. 8. They are characterized by a different dependence of the power spectrum $\omega\frac{\mathrm{d}N}{\mathrm{d}\omega}$ on $\omega$. We see that with respect to the BH regime, the other two regimes show a suppression of the gluon emission. We discuss here first the BH regime. In our MC approach, the two other regimes are implemented as a partial rejection of the BH cross section as described in the next subsection.
@@
The GB cross section $\sigma_{\text{inel}}$ depends (after integrating out energy and momentum conservation) on five kinematic variables: $\vec{k}_{{\perp}}$, the transverse momentum of the virtual gluon, $\vec{l}_{{\perp}}$, the momentum transfer from the QGP parton to the projectile and $x$, the light-cone momentum (LCM) fraction of the virtual gluon with respect to the one of the projectile. [^11] These definitions strictly apply in the cm frame where both incoming partons are colinear. It depends also on the thermal mass of the time like partons, eq. (20) or eq. (21). The propagator of the exchanged gluon is $\propto 1/(l_{T}^{2}+\mu^{2})$ with $\mu$ given by eq. (27). We perform the calculations for two extreme assumptions on the mass of the QGP partons $m_{q}$ (and $m_{g}$): Either they are set to 0 or to $\infty$, what will in the following be generically indicated by ‘‘$m_{q}=0$’’ or ‘‘$m_{q}=\infty$’’, implicitly assuming that $m_{g}=m_{q}$ for the QGP partons). The latter case will be used for calibration purposes, when comparing to the BDMPS-Z calculation, where static scattering centers are assumed.
@@
which is used in our approach. $\mathcal{M}_{\rm el}$ is the matrix element for the elastic collision, $P_{g}$ the conditional probability of emission, and $\Theta$ is the phase space constraints. The exact forms of $\mathcal{M}_{\rm el}$ and $\Theta$ depend on the choice of $m_{q}=0$ or $m_{q}\to\infty$. The exact expressions and the details of their derivation are given in eqs. (67) and (127) of appendices B and C, respectively, which also explain the MC procedure used for the sampling. The $\alpha_{s}$ in $P_{g}$ is fixed to $0.4$, whereas $\mathcal{M}_{\rm el}$ and its $\alpha_{s}$ dependence are taken as discussed in the previous subsection. The gluon spectrum in the GB approach falls off like $1/\omega$, as shown in Fig. 8,
$\omega\frac{\mathrm{d}N^{\text{GB}}}{\mathrm{d}\omega}\simeq\frac{\alpha_{s}C_{R}}{\pi}\frac{L}{\lambda_{q}}\,.$ (30)
This is the result for the massless partons, a case which is plagued by soft and colinear divergences. The generalization to finite masses – that cure the divergences – leads to some deviations with respect to this law (see eq. (105), where an extra dependence on $x$ stems from $\tilde{m}_{g}$ defined through eq. (58).
@@
  • The energy-momentum of the incoming partons is possibly corrected for the energy momentum carried by the virtual gluon and the transverse momentum it receives from the collision with the QGP parton. [^12] See discussion in subsection IV.5.2.
@@ IV.5.1 General considerations
The virtual gluons, created in inelastic collisions of the incoming parton, may scatter with the partons of the QGP. For soft gluons this has been studied by GW(P) in Gyulassy:1993hr ; Wang:1994fx and in the series of BDMPS-Z papers Baier:1996kr ; Baier:1996sk ; Zakharov:1996fv ; Zakharov:1997uu . There it has been shown that, in analogy to the LPM effect in QED, the gluon radiation amplitudes, created in subsequent scatterings of a colored projectile parton in the medium, interfere destructively. It is a distinctive new feature of QCD radiation that the virtual gluons can scatter with the partons of the QGP. These interactions modify the formation time of the gluon. As discussed, in vacuum for a gluon with four-momentum $k\equiv(\omega,\vec{k}_{\perp},k_{z})$, the formation time $t_{f}$ can be expressed as Dokshitzer:1991wu
@@
where the expression is derived assuming that $\hat{q}t_{f}\gg(k_{{\perp}})^{2}$, with $k_{{\perp}}$ being the initial transverse momentum of the radiated gluon. [^13] Since the typical $k_{\perp}\approx\mu\approx m_{g}^{\rm therm}$, this condition writes $\sqrt{\omega\hat{q}}\gg\mu^{2}$, which is satisfied provided $\omega\gg\omega_{\rm BH}$. If the parton mean free path is larger than $t_{f}$, the amplitudes for gluon emission on the trajectory of the parton do not interfere. If, however, $t_{f}$ is larger than the mean free path, interference becomes important and the relative phases of the amplitudes, $\Delta\varphi=k_{\mu}(x_{1}-x_{2})^{\mu}\approx\omega\left|\bm{x}_{1}-\bm{x}_{2}\right|(1-\cos\theta)$ have to be added. This is the LPM (BDMPS-Z) domain. For a medium of length $L$ the gluon is formed in the medium if $t_{f}\lesssim L$. The largest momentum a gluon can get on average, while traversing the medium, is $k_{\perp,\max}^{2}\simeq\hat{q}L$. We find therefore that above a limiting gluon energy of $\omega_{c}\sim\hat{q}L^{2}$, gluons can only be formed by fluctuations and are therefore strongly suppressed.
@@
In the fully coherent regime, i.e. for large gluon energy, $\omega>\omega_{c}$, the radiation spectrum is given by the first-order opacity expansion of GLV Gyulassy:2000er ; Mehtar-Tani:2019tvy  [^15] We mention that the scales, at which $\hat{q}_{g}$ should be taken in the BDMPS-Z and in the GLV regimes, differ. The first relies on a self-consistent estimation of the total transverse momentum transfer during the formation time, while the latter corresponds to $\hat{q}_{0}$, i.e. eq. (28) stripped from its Coulomb logarithm.
@@ IV.5.2 Phase accumulation algorithm
  • If the elastic collision not vetoed by the energy-momentum conservation condition, the collision counter $N_{s}$ for the virtual gluon is incremented by 1.
@@
    where $P$ and $E$ are the four-momentum and energy of the projectile which has emitted the gluon, and $k^{\prime}$ is the gluon momentum at the beginning of the time step. [^18] In practice, the time steps are small enough that this prescription is irrelevant. Hence at the end of the time step $j$ the total phase is $\sum_{l=i}^{j}\Delta\varphi_{l}$. In the limit of an energetic projectile $E\gg m_{q},\omega\gg m_{g}$, $(P\cdot k)/E$ reduces to $(m_{g}^{2}+k_{\perp}(t^{\prime})^{2})/k^{+}\approx(m_{g}^{2}+k_{\perp}(t^{\prime})^{2})/2\omega$, where $k_{\perp}(t^{\prime})$ is the instantaneous component of the virtual gluon transverse to the jet direction.
@@
  • Once a gluon is formed, it is considered as an on-shell particle which can undergo elastic and inelastic collisions. The four-momentum of the projectile parton is adapted accordingly to the outcome of the gluon formation. The two possible schemes are: a) subtract $k_{T}-l_{T}$ from the projectile parton at the moment when the virtual gluon is emitted and give it back to the projectile if and when the virtual gluon is discarded or b) subtract $k_{T}-l_{T}$ from the projectile only if and when the virtual gluon is realized. In the present study, we will place ourselves in the eikonal limit where the gluon emission leaves the momentum of the projectile parton unchanged. Therefore this correction does not play a role for the observables.
A flow diagram of the implementation of the gluon formation in our approach is presented Fig. 11. We want to mention that this algorithm is adapted to an expanding medium because it takes into account the local temperature of the QGP at which the elastic scattering of the virtual gluon takes place (and allows to stop the evolution if the QGP temperature falls below the critical temperature $T_{c}$).
@@
where the last factor represents the coherence between the vertex where the gluon is emitted from the projectile parton and the single elastic rescattering of that gluon. Due to the last factor, the radiation is suppressed if $\frac{Lk_{\perp}^{2}}{2\omega}$ is significantly larger then 0. Then
@@
For the parameter set used in the present version of our model ($\alpha_{s,{\rm eff}}$, $\kappa$,…), the value $\varphi_{c}=6$ is chosen to match the transition between the BDMPS-Z regime (eq. (34)) and the fully coherent GLV regime (eq. (40)), as documented in appendix D. In addition, a second numerical parameter is mandatory to adjust the absolute level of the $\mathrm{d}N/\mathrm{d}\omega$ MC-spectrum in the full BDPMS-Z-GLV regime. This is achieved by tuning, for large $k^{+}$, the gluon mass in the GB-seed algorithm to an effective gluon mass $m_{g}^{\rm reg}$. While the gluon mass has little influence on $\mathrm{d}N/\mathrm{d}\omega$ in the BDMPS-Z regime, being a subleading scale, it effectively affects the strength of the radiation from the GB seed — as seen in eq. (105) — and can be used as an adjustable parameter. The optimal value for $m_{g}^{\rm reg}$ is found to be less then 1% of $T$, as documented in appendix D.
@@
[…]
Gluon phase accumulation
$\varphi\to\varphi+\Delta\varphi$
[…]
Accept gluon with
probability $1/N_{s}$

### v2

@@ III.2 Medium modifications
$\hat{q}_{\rm GA}(T,p)=\hat{q}_{\rm GA}(T,p=20\text{ GeV})\times\hat{q}_{\rm cof}(p)\,,\quad\text{where}\quad\hat{q}_{\rm cof}(p)=\frac{1.69+1.25\,p}{4.07+p+0.85\ln\left(p+1\right)}\,,$ (17)
satisfying $\hat{q}_{\rm cof}(20\text{ GeV})=1$ (dimensionless). In this definition, $p$ is the dimensionless numerical value of the momentum expressed in GeV. For gluons, a color factor of $C_{A}/C_{F}$ is applied. This $p$-dependent behavior of $\hat{q}_{\rm GA}$ in Eq. (16) is similar to the model employed in  JETSCAPE:2023hqn , where $\hat{q}$ monotonously rises with the jet-parton energy. In addition, in  JETSCAPE:2023hqn coherence effects at high-virtualities are described via a $\hat{q}$ dependence on the jet-parton virtuality $Q$, which is monotonously falling. We leave the treatment of coherence effects at high virtualities for a later adaptation of our proposed algorithm and note that the approach, proposed in this section, does not depend on the explicit form of $\hat{q}$.
@@
The consequences of the in-medium induced virtuality increase is shown in Figs. 2 and 3 for initial parton energies of $10$ and $50$ GeV. The jet partons lose a large fraction of their initial virtuality without being influenced by the medium. Only later, when their virtuality is already considerably reduced and when they have traveled more than a mean free path in the medium, the contribution of the medium to their virtuality becomes visible. At this late times the medium contribution to the sum of the squared virtualities of all jet partons may reach $50\%$ of as shown on Fig. 3. For the individual tagged partons the medium contributions still may amount to more than $25\%$ of the parton virtuality at a time of $2$ to $3$ fm/c as can be seen in Fig. 2, while at later times virtualities of the tagged particles drop sharply, due to a larger amount of medium induced splittings. The aforementioned contributions depend, of course, on the choice of $\hat{q}$. A higher virtuality leads to more splittings in which the virtuality is reduced, as shown in Fig. 4. In Fig. 5 the fraction of the initial tagged partons, which have reached the minimum virtuality of $2Q_{0}$, is shown for different initial energies as a function of time. The consequences of the presence of a medium (Eq. (15) ) are also shown. The medium-induced virtuality increase (Eq. (15) has the consequence that the tagged initial quark reaches the minimum virtuality earlier due to more virtuality reducing splittings per time. Figs. 3 - 5 show that for larger jet energies the effect of the medium modifications plays a bigger role. On the one side $\hat{q}$ increases with the jet energy in our model; on the other side, initial jets with larger energies need more time to reach a low virtuality close to $2Q_{0}$. Hence the accumulated virtuality contribution from the medium ultimately becomes larger. For an initial energy $E_{\text{ini}}=10$ GeV in about $80\%$ of the jets the initial quark has reached after around $1$ fm$/c$ the minimal virtuality and has entered the low virtuality regime. For $E_{\text{ini}}=50$ GeV this happens after about $2$ fm$/c$.
@@
Figure 6: Primary Lund jet plane for the high virtuality shower for $E_{\rm ini}=50$ GeV.
Figure 7: Same as Fig 7 but with medium modifications.
In Figs. 7 and 7, we display the primary Lund plane for an initial quark-jet of $E_{\rm ini}=50\,{\rm GeV}$ evolving in the vacuum and in a medium, respectively. The upper boundary of the filled region corresponds to a gluon energy of 25 GeV and is only achievable in the first splitting. The "bulk" of the parton emission, observed for $\theta\approx 1/e$ and $k_{T}\approx 1\,{\rm GeV}/e$, corresponds to much lower values of parton energies $\approx 1\,{\rm GeV}$. The variations in the Lund plane density is due to the running of the coupling constant adopted in our model but also due to the boundaries imposed by eq. (13), which narrows the acceptable $x$-interval. On the right panel, the medium increases the overall density of emitted partons, while letting the part of the diagram located at $k_{T}\gtrsim 1\,{\rm GeV}$ unaffected. The medium-effect can be better scrutinized on Fig. 9, where the ratio of medium and vacuum emission is displayed. On this figure one distinguishes more clearly a strong enhancement of the late-time soft radiation at large angles, accompanied by a slight reduction of the soft collinear emission, resulting from the effective virtuality increase with time encoded in eq. (15).
Figure 8: The ratio between the primary Lund jet planes in Figs. 7 and 7, highlighting the effect of the medium modifications.
Figure 9: Same as Fig 9, but terminating the high virtuality according to the switch criterion expressed in eq. (3)
In Fig. 9, we show the same ratio but terminate the high virtuality according to the switch criterion of eq. (3). With this criterion and for this specific choice of parameters, the high-$Q$ stage evolves on the average down to $Q\approx 2-3\,{\rm GeV}$. Therefore, the soft collinear gluons emitted at late times in the vacuum are excluded by the switch criterion and consequently strongly depleted. For the present choice of parameters, gluons emitted in the high-$Q$ stage are mostly radiated for $t\lesssim 1\,{\rm fm}/c$, and do not suffer from the medium effect modeled with eq. (15) – see as well Fig. 2 –. This explains why the enhancement observed in Fig. 9 is not seen in Fig. 9.
@@ IV.1 Thermal masses
$m_{q}^{\text{therm}}(T)=0.087\text{ GeV}+0.7\times T\,$ (22)
with T in GeV. In the code we use these linear fit functions.
@@ IV.2 Coupling constant
During the evolution of the QGP, the elastic collisions will be probed especially at small momentum transfer. For this reason, we neglect the dependence of $\alpha_{s}$ on $Q^{2}$ and consider only an effective temperature dependence of $\alpha_{s}$ in eqs. (26) to (29). Our method relies on the models developed in ref. Gossiaux:2008jv that have led to a good agreement with experimental data in the heavy flavor (HF) sector. This is in particular true for the running model (E). We have evaluated the $\hat{q}(T,p)$ for model (E) — adapted to the light-flavor case — and then fitted $\alpha_{s}(T)$ of model (C) to reproduce at best these $\hat{q}$ values. This method leads to
@@ IV.3 Elastic collisions
Thanks to the IR regulator, the total elastic collision rate does not diverge, however its second moment - the transport coefficient $\hat{q}=\int q_{T}^{2}\frac{\mathrm{d}\Gamma_{\rm el}}{\mathrm{d}^{2}q_{T}}/\int\frac{\mathrm{d}\Gamma_{\rm el}}{\mathrm{d}^{2}q_{T}}$ does if $q_{T}$ is not limited from above. Therefore, to match the effective $\hat{q}$ of a more exact approach, model C of ref. Gossiaux:2008jv , which uses HTL propagators, we limit the transverse momentum exchange to $q_{T}^{2}<q_{T,{\rm max}}^{2}=2ET$, where $E$ is the energy of the incoming parton [^10] Not to be confused with the initial jet parton.. With this upper boundary one obtains a finite transport coefficient for jet quarks:
@@ IV.4 Inelastic collisions
The elementary inelastic collision, which we employ in our approach, is the radiation of a virtual gluon from a jet parton. The formalism, which we apply, is the QCD extension of the QED Bremsstrahlung and has been studied by Gunion and Bertsch (GB) Gunion:1981qs . As discussed, the emitted gluon with a four-vector $(\omega,k_{{\perp}},k_{z})$ needs a time $t_{f}=\omega/k_{{\perp}}^{2}$, [^11] In the numerical code, gluons are also radiated from secondary partons; in this case, the notation $\perp$” is slightly abusive, although still pretty clear. called formation time, until it becomes real. If the formation time is shorter than the mean free path $\lambda$ of the incoming parton in the medium, gluon emission from the different vertices is independent and we are in the Bethe Heitler (BH) regime. If, however, the formation time is larger than the mean free path, which depends on the temperature via the density ($\tau_{f}\propto T$), the gluon emission amplitudes interfere destructively, a process which has been studied in the BDMPS-Z approach Baier:2001yt ; Baier:1998yf ; Baier:1996sk . In this regime the gluon emission $\omega\mathrm{d}N/\mathrm{d}\omega$ is suppressed with respect to the BH regime. If the formation length exceeds the length $L$ of the QGP (assumed to be represented by a box of length L), which is the case if $\omega$ exceeds the critical value $\omega_{c}\sim\hat{q}L^{2}$, gluons will, on the average, not be formed anymore and will only be emitted by fluctuations. The three $\omega$ regions are schematically shown in Fig. 12. They are characterized by a different dependence of the power spectrum $\omega\frac{\mathrm{d}N}{\mathrm{d}\omega}$ on $\omega$. We see that with respect to the BH regime, the other two regimes show a suppression of the gluon emission. We discuss here first the BH regime. In our MC approach, the two other regimes are implemented as a partial rejection of the BH cross section as described in the next subsection.
@@
The GB cross section $\sigma_{\text{inel}}$ depends (after integrating out energy and momentum conservation) on five kinematic variables: $\vec{k}_{{\perp}}$, the transverse momentum of the virtual gluon, $\vec{l}_{{\perp}}$, the momentum transfer from the QGP parton to the projectile and $x$, the light-cone momentum (LCM) fraction of the virtual gluon with respect to the one of the projectile. [^12] These definitions strictly apply in the cm frame where both incoming partons are colinear. It depends also on the thermal mass of the time like partons, eq. (21) or eq. (22). The propagator of the exchanged gluon is $\propto 1/(l_{T}^{2}+\mu^{2})$ with $\mu$ given by eq. (28). We perform the calculations for two extreme assumptions on the mass of the QGP partons $m_{q}$ (and $m_{g}$): Either they are set to 0 or to $\infty$, what will in the following be generically indicated by ‘‘$m_{q}=0$’’ or ‘‘$m_{q}=\infty$’’, implicitly assuming that $m_{g}=m_{q}$ for the QGP partons. The latter case will be used for calibration purposes, when comparing to the BDMPS-Z calculation, where static scattering centers are assumed.
@@
which is used in our approach. $\mathcal{M}_{\rm el}$ is the matrix element for the elastic collision, $P_{g}$ the conditional probability of emission, and $\Theta$ is the phase space constraint. The exact forms of $\mathcal{M}_{\rm el}$ and $\Theta$ depend on the choice of $m_{q}=0$ or $m_{q}\to\infty$. The exact expressions and the details of their derivation are given in eqs. (73) and (135) of appendices B and B.4, respectively, which also explain the MC procedure used for the sampling. The $\alpha_{s}$ in $P_{g}$ is fixed to $0.4$, whereas $\mathcal{M}_{\rm el}$ and its $\alpha_{s}$ dependence are taken as discussed in the previous subsection. The gluon spectrum in the GB approach falls off like $1/\omega$, as shown in Fig. 12,
$\omega\frac{\mathrm{d}N^{\text{GB}}}{\mathrm{d}\omega}\simeq\frac{\alpha_{s}C_{R}}{\pi}\frac{L}{\lambda_{q}}\,$ (31)
with $\lambda_{q}$ being the mean free path. This is the result for the massless partons, a case which is plagued by soft and colinear divergences. The generalization to finite masses – that cure the divergences – leads to some deviations with respect to this law (see eq. (111), where an extra dependence on $x$ stems from $\tilde{m}_{g}$ defined through eq. (59).
@@
  • The energy-momentum of the jet parton is possibly corrected for the energy momentum carried by the virtual gluon and the transverse momentum it receives from the collision with the QGP parton. [^13] See discussion in subsection IV.5.2.
@@
[…]
Gluon phase accumulation
$\varphi\to\varphi+\Delta\varphi$
[…]
Accept gluon with
probability $1/N_{s}$
@@ IV.5.1 General considerations
The virtual gluons, created in inelastic collisions of the incoming parton, may scatter with the partons of the QGP, as shown schematically in Fig.14 .
[…]
For soft gluons this has been studied by GW(P) in Gyulassy:1993hr ; Wang:1994fx and in the series of BDMPS-Z papers Baier:1996kr ; Baier:1996sk ; Zakharov:1996fv ; Zakharov:1997uu . There it has been shown that, in analogy to the LPM effect in QED, the gluon radiation amplitudes, created in subsequent scatterings of a colored projectile parton in the medium, interfere destructively. It is a distinctive new feature of QCD radiation that the virtual gluons themselves can scatter with the partons of the QGP. These interactions modify the formation time of the gluon. As discussed, in vacuum for a gluon with four-momentum $k\equiv(\omega,\vec{k}_{\perp},k_{z})$, the formation time $t_{f}$ can be expressed as Dokshitzer:1991wu
@@
where the expression is derived assuming that $\hat{q}t_{f}\gg(k_{{\perp}}^{\rm pre})^{2}$. [^14] Since the typical $k_{\perp}\approx\mu\approx m_{g}^{\rm therm}$, this condition writes $\sqrt{\omega\hat{q}}\gg\mu^{2}$, which is satisfied provided $\omega\gg\omega_{\rm BH}$. If the parton mean free path is larger than $t_{f}$, the amplitudes for gluon emission on the trajectory of the parton do not interfere. If, however, $t_{f}$ is larger than the mean free path, interference becomes important and the relative phases of the amplitudes, $\Delta\varphi=k_{\mu}(x_{1}-x_{2})^{\mu}\approx\omega\left|\bm{x}_{1}-\bm{x}_{2}\right|(1-\cos\theta)$ have to be added. This is the LPM (BDMPS-Z) domain. For a medium of length $L$ the gluon is formed in the medium if $t_{f}\lesssim L$. The largest momentum a gluon can get on average, while traversing the medium, is $k_{\perp,\max}^{2}\simeq\hat{q}L$. We find therefore that above a limiting gluon energy of $\omega_{c}\sim\hat{q}L^{2}$, gluons can only be formed by fluctuations and are therefore strongly suppressed.
@@
In the fully coherent regime, i.e. for a large gluon energy, $\omega>\omega_{c}$, the radiation spectrum is given by the first-order opacity expansion of GLV Gyulassy:2000er ; Mehtar-Tani:2019tvy  [^16] We mention that the scales, at which $\hat{q}_{g}$ should be taken in the BDMPS-Z and in the GLV regimes, differ. The first relies on a self-consistent estimation of the total transverse momentum transfer during the formation time, while the latter corresponds to $\hat{q}_{0}$, i.e. eq. (29) stripped from its Coulomb logarithm.
@@ IV.5.2 Phase accumulation algorithm
  • If the elastic collision is not vetoed by the energy-momentum conservation condition, the collision counter $N_{s}$ for the virtual gluon is incremented by 1.
@@
    where $P$ and $E$ are the four-momentum and energy of the projectile which has emitted the gluon, and $k^{\prime}$ is the gluon momentum at the beginning of the time step. [^19] In practice, the time steps are small enough that this prescription is irrelevant. Hence at the end of the time step $j$ the total phase is $\sum_{l=i}^{j}\Delta\varphi_{l}$. In the limit of an energetic projectile $E\gg m_{q},\omega\gg m_{g}$, $(P\cdot k)/E$ reduces to $(m_{g}^{2}+k_{\perp}(t^{\prime})^{2})/k^{+}\approx(m_{g}^{2}+k_{\perp}(t^{\prime})^{2})/2\omega$ In the $E\gg m_{q},\omega\gg m_{g}$ limit the expression coincides with the one used in Zapp:2011ya . Note the extra factor 2 as compared to Zapp:2011ya , for both, for the phase accumulation formula and for the formation phase. $k_{\perp}(t^{\prime})$ is the instantaneous component of the virtual gluon transverse to the jet direction.
@@
  • Once a gluon is formed, it is considered as an on-shell particle, which can undergo elastic and inelastic collisions. The four-momentum of the projectile parton is adapted accordingly to the outcome of the gluon formation. The two possible schemes are: a) subtract $k-l$ from the projectile parton at the moment when the virtual gluon is emitted and give it back to the projectile if and when the virtual gluon is discarded or b) subtract $k-l$ from the projectile only if and when the virtual gluon is realized. In the present study, we will place ourselves in the eikonal limit where the gluon emission leaves the momentum of the projectile parton unchanged. Therefore this correction does not play a role for the observables.
A flow diagram of the implementation of the gluon formation in our approach is presented Fig. 13. The algorithm resembles the one described in Zapp:2011ya , where successive elastic scatterings of the radiation candidates increase their $k_{\perp}$, and hence their chance to become a real gluon in the medium. We want to mention that this algorithm is adapted to an expanding medium because it takes into account the local temperature of the QGP at which the elastic scattering of the virtual gluon takes place (and allows to stop the evolution if the QGP temperature falls below the critical temperature $T_{c}$).
@@
where the last factor represents the coherence between the vertex where the gluon is emitted from the projectile parton and the single elastic rescattering of that gluon. Due to the last factor, the incoherent radiation is suppressed if $\frac{Lk_{\perp}^{2}}{2\omega}$ is significantly smaller then 1.202020When $\frac{Lk_{\perp}^{2}}{2\omega}\gtrsim 1$, the cos starts to oscillate fast and averages to a subleading contribution. Then
@@
For the parameter set used in the present version of our model ($\alpha_{s,{\rm eff}}$, $\kappa$,…), the value $\varphi_{c}=6$ is chosen to match the transition between the BDMPS-Z regime (eq. (35)) and the fully coherent GLV regime (eq. (41)), as documented in appendix C. In addition, a second numerical parameter is mandatory to adjust the absolute level of the $\mathrm{d}N/\mathrm{d}\omega$ MC-spectrum in the full BDPMS-Z-GLV regime. This is achieved by tuning, for large $k^{+}$, the gluon mass in the GB-seed algorithm to an effective gluon mass $m_{g}^{\rm reg}$. While the gluon mass has little influence on $\mathrm{d}N/\mathrm{d}\omega$ in the BDMPS-Z regime, being a subleading scale, it effectively affects the strength of the radiation from the GB seed — as seen in eq. (111) — and can be used as an adjustable parameter. The optimal value for $m_{g}^{\rm reg}$ is found to be around 15% of $T$, as documented in appendix C.